# ReDel: A Toolkit for LLM-Powered Recursive Multi-Agent Systems

### v1

@@ Abstract
Recently, there has been increasing interest in using Large Language Models (LLMs) to construct complex multi-agent systems to perform tasks such as compiling literature reviews, drafting consumer reports, and planning vacations. Many tools and libraries exist for helping create such systems, however none support recursive multi-agent systems—where the models themselves flexibly decide when to delegate tasks and how to organize their delegation structure. In this work, we introduce ReDel: a toolkit for recursive multi-agent systems that supports custom tool-use, delegation schemes, event-based logging, and interactive replay in an easy-to-use web interface. We show that, using ReDel, we are able to achieve significant performance gains on agentic benchmarks and easily identify potential areas of improvements through the visualization and debugging tools. Our code, documentation, and PyPI package are open-source [^1] ReDel’s source code is available at https://github.com/zhudotexe/redel. and free to use under the MIT license.
@@ 1 Introduction
In a recursive multi-agent system, rather than a human defining the layout of multiple agents, a single root agent is given a tool to spawn additional agents. When faced with a complex task, the root agent can decompose the task into smaller subtasks, then delegate those tasks to newly-created sub-agents. Each sub-agent can then either complete the task if it is small enough, or recursively decompose and delegate the task further [^2] This is where the toolkit’s name, ReDel, comes from: it’s short for Recursive Delegation. Khot et al. (2023); Lee and Kim (2023); Prasad et al. (2024).
@@ 2 Related Work
Recent work on recursive multi-agent systems has been done by Lee and Kim (2023), Khot et al. (2023), and Prasad et al. (2024). These works introduce the method of fine-tuning or few-shot prompting LLMs to decompose complex tasks and using sub-agents to solve each part (often called recursive or hierarchical decomposition). They show that as tasks grow more complex, recursive multi-agent systems’ performance pulls away from single-agent systems. ReDel builds upon the methods introduced in these works by taking advantage of modern models’ native tool use capability Schick et al. (2023) to decompose and delegate tasks zero-shot (i.e., without human-written examples in prompt) instead of using few-shot prompting or fine-tuning. As a framework, we provide an extensible interface to apply these approaches to additional tasks and domains.
@@
Although there are other LLM-powered multi-agent system frameworks, each have various weaknesses that make them poorly suited for recursive systems and/or academic purposes. In Table 1, we compare LangGraph Campos et al. (2023), LlamaIndex Liu et al. (2022), MetaGPT Hong et al. (2024), and AutoGPT Significant Gravitas (2023) to ReDel, our system. Most are built around static multi-agent systems, with only AutoGPT supporting a single level of delegation. Only LangGraph and LlamaIndex allow agents to run in parallel asynchronously, whereas MetaGPT and AutoGPT run one agent at a time in a synchronous fashion. To log events deep within the system, only LlamaIndex provides a rigorous instrumentation suite to developers that allows them to emit events at any point while a system is running. Most do not allow developers to replay a system run from a log, with only LangGraph allowing replays by taking snapshots of each state of the system. Most do not provide a visualization interface, with only AutoGPT providing a simple chat-based UI. Unless one subscribes to a paid service, LangGraph’s replays cannot be viewed visually, and are instead presented as the raw data of each state. Finally, only AutoGPT and MetaGPT are fully open-source, with LangGraph and LlamaIndex utilizing proprietary code to offer more “premium” features beyond what their open-source libraries offer.
@@ 3.1 Tool Usage
3.2 Delegation Schemes
  • DelegateOne: Block parent agent’s execution until child agent returns its result.
  • DelegateWait: Do not block parent agent’s execution. Instead, provide a separate function to retrieve the result of a particular child.
In contrast, the DelegateWait scheme is well-suited for LLMs without parallel function calling, as it lets these models spawn multiple agents before deciding to wait on any one agent’s result. The drawback is that this runs the risk of creating zombie agents if the parent agent never retrieves the results of a particular child agent. [^3] From our testing, this is a fairly rare occurrence. As far as we are aware, ReDel is the first system to implement this type of deferred delegation scheme.
Developers can also implement their own delegation schemes modularly in a fashion similar to defining tools which can enable more complex behaviour like multi-turn delegation. We include an example of how to define a delegation scheme in Appendix A.
@@ 3.2 Delegation Schemes
{minted}
python prompt_toks = Counter() out_toks = Counter()
@@ 4.1 Experimental Setup
    WebArena: Zhou et al. (2024) Agents must do complex web tasks such as adding products to a shopping cart or commenting on GitLab.
@@ 5 Using ReDel for Error Analysis
We find that overcommitment commonly occurs when an agent performs multiple tool calls and fills its context window with retrieved information. This often, but not always, leads to failures due to the retrieved information truncating the original task from its context. In the ReDel web interface, this manifests as an abnormally small delegation graph, often consisting of only two nodes: the root node, and a single child which the root delegates to and which subsequently overcommits.

### v2

@@ Abstract
Recently, there has been increasing interest in using Large Language Models (LLMs) to construct complex multi-agent systems to perform tasks such as compiling literature reviews, drafting consumer reports, and planning vacations. Many tools and libraries exist for helping create such systems, however none support recursive multi-agent systems—where the models themselves flexibly decide when to delegate tasks and how to organize their delegation structure. In this work, we introduce ReDel: a toolkit for recursive multi-agent systems that supports custom tool-use, delegation schemes, event-based logging, and interactive replay in an easy-to-use web interface. We show that, using ReDel, we are able to easily identify potential areas of improvements through the visualization and debugging tools. Our code, documentation, and PyPI package are open-source [^1] ReDel’s source code is available at https://github.com/zhudotexe/redel. and free to use under the MIT license.
@@ 1 Introduction
In a recursive multi-agent system, rather than a human defining the layout of multiple agents, a single root agent is given a tool to spawn additional agents. When faced with a complex task, the root agent can decompose the task into smaller subtasks, then delegate those tasks to newly-created sub-agents. Each sub-agent can then either complete the task if it is small enough, or recursively decompose and delegate the task further [^2] This is where the toolkit’s name, ReDel, comes from: it’s short for Recursive Delegation. Khot et al. (2023); Lee and Kim (2023); Prasad et al. (2024) (Figure 1).
@@ 2 Related Work
Recent work on recursive multi-agent systems has been done by Lee and Kim (2023), Khot et al. (2023), Qi et al. (2023), and Prasad et al. (2024). These works introduce the method of fine-tuning or few-shot prompting LLMs to decompose complex tasks and using sub-agents to solve each part (often called recursive or hierarchical decomposition). ReDel builds upon the methods introduced in these works by taking advantage of modern models’ native tool use capability Schick et al. (2023) to decompose and delegate tasks zero-shot (i.e., without human-written examples in prompt) instead of using few-shot prompting or fine-tuning. As a framework, we provide an extensible interface to apply these approaches to additional tasks and domains.
Other multi-agent system methods such as agent evolution Qian et al. (2024); Yuan et al. (2024); Zhou et al. (2024b) perturb human-written prompts and tools to create new variations of sub-agents on the fly. In this paper, we choose to explore delegation using zero-shot prompting and function calling without on-the-fly adaptation, but our framework is flexible enough to implement these alternate methods of agent delegation as well.
@@
Although there are other LLM-powered multi-agent system frameworks, each have various weaknesses that make them poorly suited for recursive systems and/or academic purposes. In Table 1, we compare LangGraph Campos et al. (2023), LlamaIndex Liu et al. (2022), MetaGPT Hong et al. (2024), AutoGPT Significant Gravitas (2023), and XAgent XAgent Team (2023) to ReDel, our system. Most are built around static multi-agent systems, with only AutoGPT and XAgent supporting a single level of delegation. Only LangGraph and LlamaIndex allow agents to run in parallel asynchronously, whereas MetaGPT, AutoGPT, and XAgent run one agent at a time in a synchronous fashion. To log events deep within the system, only LlamaIndex provides a rigorous instrumentation suite to developers that allows them to emit events at any point while a system is running. Most do not allow developers to replay a system run from a log, with only LangGraph allowing replays by taking snapshots of each state of the system. Most do not provide a visualization interface, with only AutoGPT and XAgent providing a simple chat-based UI. Unless one subscribes to a paid service, LangGraph’s replays cannot be viewed visually, and are instead presented as the raw data of each state. Finally, only AutoGPT, MetaGPT, and XAgent are fully open-source, with LangGraph and LlamaIndex utilizing proprietary code to offer more “premium” features beyond what their open-source libraries offer.
@@ 3.1 Tool Usage
python prompt_toks = Counter() out_toks = Counter()
@@
3.2 Delegation Schemes
@@
  • DelegateOne: Synchronously block the parent agent’s execution until the child agent returns its result (in the form of its chat output).
  • DelegateWait [^3] Named so in that it provides two functions to agents: delegate(), which sends the instructions to the child agent and spawns it, and wait(), which retrieves its result, waiting for it to finish if necessary.: Do not block parent agent’s execution. Instead, provide a separate function to asynchronously retrieve the result (chat output) of a particular child.
In contrast, the DelegateWait scheme is well-suited for LLMs without parallel function calling, as it lets these models spawn multiple agents before deciding to wait on any one agent’s result (i.e., retrieve its conversational output). The drawback is that this runs the risk of creating zombie agents if the parent agent never retrieves the results of a particular child agent. [^4] From our testing, this is a fairly rare occurrence. As far as we are aware, ReDel is the first system to implement this type of deferred delegation scheme.
Developers can also implement their own delegation schemes modularly in a fashion similar to defining tools which can enable more complex behaviour. For example, a developer might implement a delegation scheme that allows a parent agent to ask follow-up questions to existing children to enable multi-turn delegation. Developers can also use the delegation scheme to control how the child passes information back to its parent – for example, having each child call a set_result() function to explicitly record its answer to a subtask instead of implicitly sending its chat output to the parent. We include examples of how to define a delegation scheme in Appendix A and in our GitHub repository.
@@ 4.1 Experimental Setup
    WebArena: Zhou et al. (2024a) Agents must do complex web tasks such as adding products to a shopping cart or commenting on GitLab.
@@ 5 Using ReDel for Error Analysis
We find that overcommitment commonly occurs when an agent performs multiple tool calls and fills its context window with retrieved information. In the ReDel web interface, this manifests as an abnormally small delegation graph, often consisting of only two nodes: the root node, and a single child which the root delegates to and which subsequently overcommits. In practice, this often, but not always, results in the overcommitting model “forgetting” the task it was meant to accomplish due to the original task being truncated its limited context window. An overcommitting model might fail a task because it outputs a summary of whatever remains in its context window instead of the answer to the original task, whereas a task failure due to causes other than overcommittment might look like a hallucinated result or a simple apology for being unable to complete the task.
@@ 6 Conclusion
Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship, under Grant No. DGE-2236662. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or views, either expressed or implied, of ODNI, IARPA, the NSF, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.